OpenStudy (anonymous):

Write the complex number in polar form. Express the argument in degrees. -3 + 4i

OpenStudy (kc_kennylau):

Polar form is writing a complex number in the form of \(r(\cos\theta+i\sin\theta)\), where \(r=\sqrt{a^2+b^2}\) and \(\theta=\tan^{-1}\frac ba\)